# “I Want It That Way”: Enabling Interactive Decision Support Using Large Language Models and Constraint Programming

### v1

@@ Abstract.
A critical factor in the success of decision support systems is the accurate modeling of user preferences. Psychology research has demonstrated that users often develop their preferences during the elicitation process, highlighting the pivotal role of system-user interaction in developing personalized systems. This paper introduces a novel approach, combining Large Language Models (LLMs) with Constraint Programming to facilitate interactive decision support. We study this hybrid framework through the lens of meeting scheduling, a time-consuming daily activity faced by a multitude of information workers. We conduct three studies to evaluate the novel framework, including a diary study (n=64) to characterize contextual scheduling preferences, a quantitative evaluation of the system’s performance, and a user study (n=10) with a prototype system. Our work highlights the potential for a hybrid LLM and optimization approach for iterative preference elicitation and design considerations for building systems that support human-system collaborative decision-making processes.
@@ 1. Introduction
Understanding what a user likes and dislikes, i.e., their preferences, is crucial for the success of intelligent systems that provide decision and negotiation support across many domains. Decision support systems target the decision-making process itself, helping the user discover and articulate their preferences, understand the link between preferences and decision outcomes, and analyze the steps taken in the process. However, eliciting preference information from a user to inform such a decision process is challenging. This challenge manifests not only as a limitation of the algorithms for elicitating or representing preferences but also as an inherent difficulty of users knowing about or expressing their desired solutions.
@@
Individuals are often unsure of their preferences, hold inconsistent preferences, and change their preferences as they go about making decisions which can be attributed to the inherent limitations of individuals’ memory, attention, and knowledge  (Lichtenstein and Slovic, 2006; Slovic, 1995; Payne et al., 1992, 1993). These challenges are shared across widespread decision-making scenarios, such as meeting scheduling, travel planning or purchasing a house. Preference elicitation tools and techniques (Goldsmith and Junker, 2008) are known to assist human decision-makers in overcoming some of their inherent memory or knowledge limitations, extracting information from the user’s mental representation of that preference and translating it into a representation through which the system can reason. However, the constructive nature of human preferences implies that the interaction between the system and a user can greatly influence the quality of the preference information (Chen and Pu, 2009; Pommeranz et al., 2012; Johnson et al., 2005) and the user’s acceptance of the results provided by the system (Carenini and Poole, 2002; Bettman et al., 1998). Prior work has shown that such preference elicitation systems need to be collaborative to facilitate seamless information exchange between the user and the system  (Peintner et al., 2008), contextualized to the decision-making process (Payne et al., 1993; Slovic, 1995; Pu et al., 2003; Carenini and Poole, 2002), and incremental to account for evolving preferences (Pu et al., 2003; Pommeranz et al., 2012; Ardissono et al., 2003).
@@
However even once preferences are elicited, finding a solution that best meets them represents a challenging decision making task. Optimization models such as Constraint Programming (CP) are a natural fit for solving this complex reasoning problem and have been successfully applied in a range of multi-attribute decision problems. CP specifically aligns with multi-attribute decision theory (Edwards and Newman, 1982; Keeney and Raiffa, 1993; Hammond et al., 2015) where preferences are represented as utility functions or embedded as constraints, and constraint satisfaction techniques are leveraged to manage trade-offs  (Simon et al., 2004). However, traditional optimization approaches require full knowledge of the model’s specification—including a fully specified objective and all relevant constraints. This is at odds with how users approach complex and contextual decision-making tasks where they often do not know relevant preferences unless presented with candidate solutions. Furthermore, customizing optimization models requires expertise in coding and optimization, an unrealistic barrier for most users. Finally, these optimization approaches must function as dynamic and incremental reasoning engines: users actively construct and adjust their preferences while the system continuously learns and embeds their preferences, offering increasingly improved suggestions until a satisfactory one is achieved.
@@
In this paper, we explore an innovative framework that supports iterative preference elicitation, aligning with how humans develop or construct preferences in their decision-making. Elicited preferences are embedded into specific constraints that can be solved within an optimization setting. We leverage (1) Large Language Models (LLMs) to facilitate natural and intuitive communication of preferences and suggestions as well as to reason about how preferences should be incorporated and translated into structured constraint functions and (2) CP techniques to reason over the space of outcomes and find optimal solutions given incorporated constraints.
@@
We evaluate this framework for interactive decision support systems through the lens of meeting scheduling—a routine, tedious, and time-consuming daily activity faced by a multitude of information workers. At its core, scheduling a meeting is a complex spatial-temporal task that is highly contextual, and finding a suitable meeting time may require reasoning over conflicting preferences, constraints, and priorities (for double-blind review, 2023). It is also an incremental and continuous process that requires iteratively refining candidate solutions in a more or less greedy fashion (Brzozowski et al., 2006). Therefore, interactive decision support tools are a natural fit for scheduling as they can provide contextualized and interactive support for the complex reasoning that is necessary to complete the task.
@@
We introduce MeetMate, a hybrid LLM and CP approach that capitalizes on the benefits of both technologies to build an interactive decision support system for meeting scheduling. MeetMate  is designed to facilitate human-system collaboration that encompasses two key components: Preference Construction and Preference Incorporation. Preference Construction is the process by which users of the system evaluate candidate decisions (i.e., suggested meeting times) and express new preferences and constraints that need to be integrated into the system. Preference Incorporation is a mirror of the construction process from the system perspective and involves embedding expressed user preferences in a form that can be used by the system to generate new suggested decisions. We refer to the entire process of iteratively eliciting preferences via generating and critiquing suggestions as Preference Elicitation. Because preference elicitation is an iterative process, MeetMate  unifies Preference Construction and Preference Incorporation into one dynamic loop where preferences are elicited in conjunction with refined suggestion generations. Figure 1 outlines the key components of our interactive decision support system for meeting scheduling.
@@
We conducted three studies to better understand preference elicitation in the context of meeting scheduling. First, we study how preferences are constructed via a diary study that collected naturalistic preferences in the moment of scheduling (n=64, Section 3). Informed by the results of the diary study, we then introduce MeetMate (Section 4) and quantitatively evaluate the system’s capabilities for preference incorporation (Section 5). Finally, we conduct a user study (n=10, Section 6) with a prototype system to observe the iterative preference elicitation flow in situ and to inform future design improvements for interactive decision support systems. Through our study, we confirm that participants’ preferences are indeed constructed through the iterative preference elicitation process of evaluating dynamic time suggestions that incorporate their feedback. We also identify that system explanations play a huge role in bridging the gap between understanding the system and expressing preferences, and contrasted suggestions not only give users a sense of agency but also help make the decision-making process efficient. Our work highlights the potential for a hybrid LLM and optimization approach for iterative preference elicitation and design considerations for building systems that support human-system collaborative decision-making processes.
@@
Figure 1. An overview of the interactive loop for conversational decision support. The loop is initiated via an initial time suggestion (or set of suggestions) by the system presented to the user (A). During Preference Construction users evaluate a proposed suggestion (C) and either express a new preference (D) to improve the given suggestion, or accept the current suggestion. New preferences are integrated into the system during Preference Incorporation which requires the system to both embed the stated preference into the system (W) and then use it to generate a new time suggestion (F). This process is iterated until a suitable time is found.
@@ 2.1. Human Preferences and Design Implications
In 2002, Carenini and Poole (2002) described a shift in classical decision theory towards constructive preferences, contrary to the prevailing view of stable and rational preferences (e.g., (Doyle, 2004)). Psychological studies have shown that preferences are often constructed rather than stable, and individuals may hold inconsistent or evolving preferences (Slovic, 1995; Lichtenstein and Slovic, 2006; Johnson et al., 2005). This constructive nature of preferences has been emphasized, suggesting that preferences are more like architectural creations than archaeological discoveries (Gregory et al., 1993; Payne et al., 2000). In various decision-making theories (Janis and Mann, 1977; Mills and O’neal, 1971; Montgomery and Willén, 1999; Russo et al., 1996), it has been posited that people construct preferences when faced with decisional conflicts. Preferences are not well-defined in most situations but are constructed in the context of decision-making, particularly when conflicts arise. This process allows individuals to adjust their initial preferences to make confident decisions, especially when one option clearly outweighs others (Simon et al., 2008). Incremental preference elicitation has received attention (Pu et al., 2003; Pommeranz et al., 2012; Ardissono et al., 2003) due to the uncertain nature of user goals and preferences. It involves improving a user’s preference model over time as they interact with the system.
@@
Preference construction has also been explored within the framework of multi-attribute decision theory (MAUT) (Edwards and Newman, 1982; Keeney and Raiffa, 1993; Hammond et al., 2015), where preferences are represented as utility functions. Research in this area has shown that people tend to increase their preferences for attributes of chosen options while decreasing preferences for attributes of rejected options. Similar effects have been observed in negotiation settings (Curhan et al., 2004). These findings align with the goal of maximizing the ease of justifying a decision, as identified by Bettman et al. (1998). This research on decision-making by constraint satisfaction has identified coherence shifts in various tasks, including high-level reasoning, social reasoning, and factual inferences (Simon et al., 2004).
@@
To develop an effective decision support system, designers must take into account the process of human preference development (Pu et al., 2003; Pommeranz et al., 2012; Chen and Pu, 2004). Relevant literature provides key insights for designing effective decision support systems in two main areas. First, a wealth of research in HCI and user-centered design highlights the paramount significance of interactivity and collaboration within decision support systems. Effective decision support necessitates a collaborative environment wherein both the user and the system can seamlessly exchange information and jointly tackle decision-making challenges. Second, studies in supporting preference construction offer further guidance on structuring this collaborative process. This collaboration involves two critical elements: (1) natural interaction, i.e., the system UI should offer user-friendly tools for users to express, revise, and provide feedback on their preferences and constraints, as well as critique system suggestions; (2) iterative and transparent system recommendations, i.e., the system should contextualize the decision-making, enabling users to understand decision consequences (Payne et al., 1993; Slovic, 1995; Pu et al., 2003). Providing concrete examples for user critique (Pu et al., 2003; Chen and Pu, 2012; McGinty and Reilly, 2010), along with immediate feedback and explanations, helps users grasp the correlation between preferences, constraints, and proposed solutions (Carenini and Poole, 2002).
@@
Our work primarily centers on the information flow from the system to the user, aligning with the process of human preference development and decision-making. It underscores the significance of contextualized elicitation mechanisms, concrete examples for user critique, immediate feedback with explanations, incremental suggestion generation, and the utilization of constraint satisfaction as the reasoning algorithm.
@@ 2.2. Preference Elicitation Techniques
To create a user-friendly system, designers must consider the development of human preferences, and as such, preference elicitation is a well-explored topic in artificial intelligence, recommender systems, and human-computer interaction (Goldsmith and Junker, 2008). Various techniques exist for inferring user preferences, including rankings, corrections, critiques, ratings, queries, and past behavior. Preference elicitation methods can be broadly categorized into two types: implicit and explicit. Implicit methods involve learning user preferences from behavior, aligning with Revealed Preference Theory (Richter, 1966). In contrast, explicit methods, such as collecting ratings, actively involve users in constructing and reflecting on their preferences. Our approach falls into the explicit category, allowing users to specify their preferences during the process. Our method draws inspiration from techniques like Similarity and Tweaking (Carenini et al., 2003; Jannach et al., 2021); and Example-Critiquing Interaction (Chen and Pu, 2012). These approaches involve users in the recommendation process by having them specify attributes or critique candidate options. Similarly, our system follows an iterative approach, enabling users to critique time recommendations within the chat interface.
@@
Our work is also related to preference elicitation in Decision Support Systems (DSS), where interactive tools assist users in decision-making. Most DSS rely on multi-attribute utility theory (Keeney and Raiffa, 1993) to represent preferences as utility functions. However, constructing these utility functions in collaboration with users is challenging. Traditional methods like absolute measurement and pairwise comparison (active elicitation) (Aloysius et al., 2006) may lead to misalignment between users’ qualitative mental models of preferences and the system’s quantitative representation. In the context of calendar scheduling, understanding user preferences is essential. Previous research has used various approaches, including machine learning on existing calendars (Mitchell et al., 1994), showing users sample schedules (Brzozowski et al., 2006), and allowing direct input of parameters (Haynes et al., 1997). Our approach stands out by allowing users to specify new preferences flexibly, accommodating more diverse contextual scheduling preferences.

### v3

@@ Abstract.
A critical factor in the success of many decision support systems is the accurate modeling of user preferences. Psychology research has demonstrated that users often develop their preferences during the elicitation process, highlighting the pivotal role of system-user interaction in developing personalized systems. This paper introduces a novel approach, combining Large Language Models (LLMs) with Constraint Programming to facilitate interactive decision support. We study this hybrid framework through the lens of meeting scheduling, a time-consuming daily activity faced by a multitude of information workers. We conduct three studies to evaluate the novel framework, including a diary study to characterize contextual scheduling preferences, a quantitative evaluation of the system’s performance, and a user study to elicit insights with a technology probe that encapsulates our framework. Our work highlights the potential for a hybrid LLM and optimization approach for iterative preference elicitation, and suggests design considerations for building systems that support human-system collaborative decision-making processes.
Decision Support, Large Language Models, Constraint Programming, Preference Elicitation, Meeting Scheduling
copyright: acmlicensedjournal: TIISjournalyear: 2024journalvolume: 14journalnumber: 3article: 22publicationmonth: 9doi: 10.1145/3685053ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Understanding what a user likes and dislikes, that is, their preferences, is often crucial for the success of intelligent systems that provide decision and negotiation support across many domains. Decision support systems assist the decision-making process itself, helping the user discover and articulate their preferences, understand the link between preferences and decision outcomes, and analyze the steps taken in the process. However, eliciting preference information from a user to inform such a decision process is challenging. This challenge manifests not only as a limitation of the algorithms for elicitating or representing preferences but also as an inherent difficulty of users knowing about or expressing their desired solutions.
@@
A substantial body of literature on judgment and decision-making suggests that preferences are constructed at the time of decision-making rather than being pre-determined or stable (Lichtenstein and Slovic, 2006; Slovic, 1995). Individuals are often unsure of their preferences, hold inconsistent preferences, and change their preferences as they go about making decisions, which can be attributed to the inherent limitations of individuals’ memory, attention, and knowledge  (Lichtenstein and Slovic, 2006; Slovic, 1995; Payne et al., 1992, 1993). These challenges are shared across widespread decision-making scenarios, such as meeting scheduling, travel planning or purchasing a house. Preference elicitation tools and techniques (Goldsmith and Junker, 2008) are used to assist human decision-makers in overcoming some of their inherent memory or knowledge limitations, extracting information from the user’s mental representation of that preference and translating it into a representation through which the system can reason. However, the constructive nature of human preferences implies that the interaction between the system and a user can greatly influence the quality of the preference information (Chen and Pu, 2009; Pommeranz et al., 2012; Johnson et al., 2005) and the user’s acceptance of the results provided by the system (Carenini and Poole, 2002; Bettman et al., 1998). Prior work has shown that such preference elicitation systems need to be (1) collaborative to facilitate seamless information exchange between the user and the system (Peintner et al., 2008), (2) contextualized to the decision-making process (Payne et al., 1993; Slovic, 1995; Pu et al., 2003; Carenini and Poole, 2002), and (3) incremental to account for evolving preferences (Pu et al., 2003; Pommeranz et al., 2012; Ardissono et al., 2003). In building a preference elicitation system, Large Language Models (LLMs) provide a promising opportunity to support the naturalistic and contextual elicitation of user preferences via chat. Recent work has demonstrated their potential to work on collaborative tasks (Lin et al., 2024), and its integration into chat enables a system that can be responsive to both the context of the decision-making task and the incremental evolution of preferences.
@@
Even once preferences are elicited, finding a solution that best meets them represents a challenging decision-making task. Optimization techniques such as Constraint Programming (CP) are a natural fit for solving this complex reasoning problem and have been successfully applied in a range of multi-attribute decision problems. One major shortcoming of traditional optimization approaches is that they require full knowledge of the model’s specification—including a fully specified objective and all relevant constraints in a structured format. This is at odds with how users approach complex and contextual decision-making tasks where they often do not know relevant preferences unless presented with candidate solutions. Furthermore, customizing optimization models requires expertise in coding and optimization, an unrealistic barrier for most users. Here, in addition to their potential in eliciting preferences, LLMs provide another opportunity to overcome this technical hurdle by translating natural language into structurued forms that can be used by an underlying optimization solver. Finally, these optimization approaches must function as dynamic and incremental reasoning engines: users actively construct and adjust their preferences while the system continuously learns and embeds their preferences, offering increasingly improved suggestions until a satisfactory one is achieved.
@@
In this paper, we explore a framework that supports iterative preference elicitation, aligning with how humans develop or construct preferences in their decision-making. Elicited preferences are embedded into specific constraints that can be solved within an optimization setting. We leverage (1) LLMs to facilitate natural and intuitive communication of preferences and suggestions as well as to reason about how preferences should be incorporated and translated into structured constraint functions, and (2) CP techniques to reason over the space of outcomes and find optimal solutions given incorporated constraints.
@@
We study this framework for interactive decision support through the lens of meeting scheduling—a routine, tedious, and time-consuming daily activity faced by a multitude of information workers. At its core, scheduling a meeting is a complex spatio-temporal task that is highly contextual, and finding a suitable meeting time may require reasoning over conflicting preferences, constraints, and priorities (Sun et al., 2024). It is also an incremental and continuous process that requires iteratively refining candidate solutions in a more or less greedy fashion (Brzozowski et al., 2006). Therefore, interactive decision support tools are a natural fit for scheduling as they can provide contextualized and interactive support for the complex reasoning that is necessary to complete the task. We specifically study professional meeting scheduling for information workers. This setting comes with a host of implicit scheduling constraints including working hours, implicit power dynamics (i.e., for managers vs. team members), and imperfect calendar information (i.e., undisclosed personal scheduling constraints that are not reflected in a work calendar) that provide a rich setting in which to study contextual preference elicitation. Moreover, our work focuses on the process of a meeting organizer finding a candidate meeting time. This is in line with most modern industry-scale tools for meeting scheduling, such as Microsoft Outlook, where attendees’ information is limited to their calendar’s free/busy time. Previous work has also shown that organizers typically have more power than other participants over where a meeting lands (Mok et al., 2023). We fully acknowledge that meeting scheduling is a distributed negotiation task amongst multiple participants and encourage follow-up research that builds on our organizer-centric investigation.
@@
We introduce MeetMate, a hybrid LLM and CP approach that capitalizes on the benefits of both technologies to build an interactive decision support system for meeting scheduling. MeetMate  is designed to facilitate human-system collaboration that encompasses two key components: Preference Construction and Preference Incorporation. Preference Construction is the process by which users of the system evaluate candidate decisions (i.e., suggested meeting times) and express new preferences and constraints that need to be integrated into the system. Preference Incorporation is a mirror of the construction process from the system perspective and involves embedding expressed user preferences in a form that can be used by the system to generate new suggested decisions. We refer to the entire process of iteratively eliciting preferences via generating and critiquing suggestions as Preference Elicitation. Because eliciting preferences is an iterative process, MeetMate  unifies Preference Construction and Preference Incorporation into one dynamic loop where preferences are elicited in conjunction with refined suggestion generations. Figure 1 outlines the key components of our interactive decision support system for meeting scheduling.
@@
We conducted three studies to understand preference elicitation in the context of meeting scheduling. First, we study how preferences are constructed via a diary study that collected naturalistic preferences in the moment of scheduling (Section 3). Informed by the results of the diary study, we then introduce MeetMate (Section 4) and quantitatively evaluate the system’s capabilities for preference incorporation (Section 5). Our results show that LLMs can generate high-quality functional representations of scheduling constraints that achieve high precision and recall when compared to human expert implementations of the constraints. Finally, we use MeetMate  as a technology probe to conduct a user study (Section 6) , to observe the iterative preference elicitation flow in situ, and to inform future design improvements for intelligent interactive decision support systems. Through our study, we confirm that participants’ preferences are indeed constructed through the iterative preference elicitation process of evaluating dynamic time suggestions that incorporate their feedback. We also identify that system explanations play an important role in bridging the gap between understanding the system and expressing preferences, and contrasted suggestions not only give users a sense of agency but also help make the decision-making process efficient. Our work highlights the potential for a hybrid LLM and optimization approach for iterative preference elicitation and design considerations for building systems that support human-system collaborative decision-making processes.
@@
Figure 1. An overview of the interactive loop for conversational decision support. The loop is initiated via an initial time suggestion (or set of suggestions) by the system presented to the user (A). During Preference Construction users evaluate a proposed suggestion (C) and either express a new preference (D) to improve the given suggestion, or accept the current suggestion. New preferences are integrated into the system during Preference Incorporation which requires the system to both embed the stated preference into the system (E) and then use it to generate a new time suggestion (F). This process is iterated until a suitable time is found.
[…]
Two components: Preference Construction and Preference Incorporation. Preference Construction is driven by the User and consists of (B) Suggestion Evaluation followed by (C) Preference Expression, which passes a (D) Preference to the System. Preference Incorporation is driven by the System and contains (E) Preference Embedding followed by (F) Suggestion Generation, which passes a (A) Time Suggestion to the User.
@@ 2.1. Human Preferences and Design Implications
In 2002, Carenini and Poole (2002) described a shift in classical decision theory towards constructive preferences, contrary to the prevailing view of stable and rational preferences (e.g., (Doyle, 2004)). Psychological studies have shown that preferences are indeed often constructed rather than stable, and individuals may hold inconsistent or evolving preferences (Slovic, 1995; Lichtenstein and Slovic, 2006; Johnson et al., 2005). This constructive nature of preferences has been emphasized, suggesting that preferences are more like architectural creations than archaeological discoveries (Gregory et al., 1993; Payne et al., 2000). In various decision-making theories (Janis and Mann, 1977; Mills and O’Neal, 1971; Montgomery and Willén, 2007; Russo et al., 1996), it has been posited that people construct preferences when faced with decisional conflicts. Preferences are not well-defined in most situations but are constructed in the context of decision-making, particularly when conflicts arise. This process allows individuals to adjust their initial preferences to make confident decisions, especially when one option clearly outweighs others (Simon et al., 2008). Incremental preference elicitation has received attention (Pu et al., 2003; Pommeranz et al., 2012; Ardissono et al., 2003) due to the uncertain nature of user goals and preferences. It involves improving a user’s preference model over time as they interact with a system.
@@
Preference construction has also been explored within the framework of multi-attribute utility theory (MAUT) (Edwards and Newman, 1982; Keeney and Raiffa, 1993; Hammond et al., 2015), where preferences are represented as utility functions. Research in this area has shown that people tend to increase their preferences for attributes of chosen options while decreasing preferences for attributes of rejected options. Similar effects have been observed in negotiation settings (Curhan et al., 2004). These findings align with the goal of maximizing the ease of justifying a decision, as identified by Bettman et al. (1998).
@@
To develop an effective decision support system, designers must take into account the process of human preference development (Pu et al., 2003; Pommeranz et al., 2012; Chen and Pu, 2004). Relevant literature provides key insights for designing effective decision support systems in two main areas. First, research in human-computer interaction (HCI) and user-centered design highlights the paramount significance of interactivity and collaboration within decision support systems. Effective decision support necessitates a collaborative environment wherein both the user and the system can seamlessly exchange information and jointly tackle decision-making challenges. Second, studies in supporting preference construction offer further guidance on structuring this collaborative process. The collaboration involves two critical elements: (1) natural interaction, that is, the system user interface (UI) should offer user-friendly tools for users to express, revise, and provide feedback on their preferences and constraints, as well as critique system suggestions; (2) iterative and transparent system recommendations, that is, the system should contextualize the decision-making, enabling users to understand decision consequences (Payne et al., 1993; Slovic, 1995; Pu et al., 2003). Providing concrete examples for user critique (Pu et al., 2003; Chen and Pu, 2012; McGinty and Reilly, 2010), along with immediate feedback and explanations, helps users grasp the correlation between preferences, constraints, and proposed solutions (Carenini and Poole, 2002).
@@
Our work primarily centers on the information flow from the system to the user, aligning with the process of human preference development and decision-making. It underscores the significance of contextualized elicitation mechanisms, concrete examples for user critique, immediate feedback with explanations, incremental suggestion generation, and the utilization of constraint satisfaction as part of the reasoning engine.
@@ 2.2. Preference Elicitation Techniques
Preference elicitation has been studied in artificial intelligence (AI), recommender systems, and HCI (Goldsmith and Junker, 2008). Various techniques exist for inferring user preferences, including rankings, corrections, critiques, ratings, queries, and past behavior. Preference elicitation methods can be broadly categorized into two types: implicit and explicit. Implicit methods involve learning user preferences from behavior, aligning with Revealed Preference Theory (Richter, 1966). In contrast, explicit methods, such as collecting ratings, actively involve users in constructing and reflecting on their preferences. Our approach falls into the explicit category, allowing users to specify their preferences during the decision-making process. Our method draws inspiration from techniques like Similarity and Tweaking (Carenini et al., 2003; Jannach et al., 2021), and Example-Critiquing Interaction (Chen and Pu, 2012). These approaches involve users in the recommendation process by having them specify attributes or critique candidate options. Similarly, our system follows an iterative approach, enabling users to critique time recommendations within a chat interface.
@@
Our work is also related to preference elicitation in decision support systems, where interactive tools assist users in decision-making. Many decision support systems rely on MAUT (Keeney and Raiffa, 1993) to represent preferences as utility functions. However, constructing these utility functions in collaboration with users is challenging. Traditional methods like absolute measurement and pairwise comparison (active elicitation) (Aloysius et al., 2006) may lead to misalignment between users’ qualitative mental models of preferences and the system’s quantitative representation.
2.3. User Scheduling Preferences
Whether a user is arranging a new meeting or responding to a request, they consider multiple objectives, such as minimizing disruptions, travel time, and preferred locations (Berry et al., 2007; Oh and Smith, 2005). An effective calendar scheduling system needs to understand these user preferences to provide useful recommendations (Oh and Smith, 2005), which has inspired a line of work on eliciting and incorporating user preferences into calendar scheduling systems (Dent et al., 1992; Mitchell et al., 1994; Sen et al., 1997; Chajewska et al., 2000; Tullio et al., 2002; Oh and Smith, 2005; Gervasio et al., 2005; Brzozowski et al., 2006; Krzywicki et al., 2010; Kim et al., 2018). Prior work has explored complex and often contradictory preferences in the context of meeting scheduling (see (Sun et al., 2024) for a broader review). Understanding user scheduling preferences is a complicated task that suffers from many of the same challenges as the broader preference elicitation setting. Users may not be fully aware of their preferences, and may find the elicitation process itself burdensome. Attempting to address the burden of explicit preference elicitation, prior work has used machine learning on user’s existing calendar to learn scheduling preferences (Mitchell et al., 1994; Brzozowski et al., 2006; Gervasio et al., 2005). There have also been works focused on explicit preference elicitation by showing users sample schedules (Brzozowski et al., 2006; Viappiani et al., 2006), or allowing them to directly input parameters into a fixed preference model (Haynes et al., 1997). Berry et al. (2007) took a user-centered approach, capturing criteria through structured interviews and using an interactive interface. However, all these approaches require a fixed preference model that users, either implicitly or explicitly, populate with their own preferences. In contrast, our approach allows users to specify new preferences that may not be captured in a fixed preference model (e.g., realizing they want a 15 minute break only after seeing more than 2 hours of back-to-back meetings). This flexibility better aligns with the diversity of contextual scheduling preferences we captured in our diary study (Section 3).
Given the routine, tediuous, and time-consuming nature of meeting scheduling, a wide array of tools have been developed to help users schedule meetings (Brzozowski et al., 2006; Cranshaw et al., 2017; Kim et al., 2018; Oh and Smith, 2005; Berry et al., 2011). However, many existing calendar tools are too rigid to accommodate the diversity, uncertainty, and evolution of user scheduling preferences. Recent work has tried to automate the task of finding a candidate meeting time via machine learning (Brzozowski et al., 2006), but gives users no recourse to customize or tweak suggested times if the ones provided do not work. Other tools, such as Doodle and When2meet, have focused on the group nature of meeting scheduling, allowing users to share their availability and find a shared meeting time. However, these tools typically only allow users to give binary feedback (i.e., whether or not they are available for a given time slot) and do not allow users to negotiate over subjective preferences and their relative importances. For instance, users may not want to put morning availability for a meeting unless it is the only option for which all participants are available. In contrast, our proposed system allows users to specify scheduling preferences and their relative importance in natural language and can accommodate more diverse preferences then existing systems.